Identify an equation in standard form for ellipse with the center at the origin, vertex at (3, 0), and focus at (1, 0).

Answer by lwsshak3(11628)

Identify an equation in standard form for ellipse with the center at the origin, vertex at (3, 0), and focus at (1, 0).
**
Standard form of equation for an ellipse with horizontal major axis:
(x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
For given ellipse:
Center: (0,0)
length of horizontal major axis=6=2a (between end points of vertices)
a=3
a^2=9
c=1(from ctr to focal point)
c^2=1
c^2=a^2-b^2
b^2=a^2-c^2=9-1=8
Equation:
x^2/9+y^2/8=1